Presentation is loading. Please wait.

Presentation is loading. Please wait.

MICE - UK Project Progress Roy Preece 30 th April 2013.

Similar presentations

Presentation on theme: "MICE - UK Project Progress Roy Preece 30 th April 2013."— Presentation transcript:

1 MICE - UK Project Progress Roy Preece 30 th April 2013

2 Overview Focus Coil Magnet 1 Cooling / Testing Magnet 2 Construction Progress Magnetic Mitigation Analysis West Wall Rack Room 2 and Control room Partial return yoke RF DL and Hall operations Hall operations Decay Solenoid Spectrometer Solenoid frames LH2

3 Focus Coil – Magnet 1 Cooling / Testing Test of the cryo coolers – 23 rd November –Successful operation with first and second stage cooling. –Water chiller operated well –Confidence for starting the LN2 cooling LN2 cooling – 26 th November –Goal of 90K for the cold mass –LN2 cooling during the day –Pump and purge with He gas 3 times –Cryo cooler operate overnight. –Allow second stage of cold head to warm to 80K before re-starting LN2 input. –Final LN2 input 6 th December – Cold mass at 90K Coolers operating – 6 th December –He gas regulated to input 1.05 bar to the cold mass Helium circuit –Approximately 20 bar of gas has been used (not lost) –Cold mass and second stage heads at temperatures that indicate Helium being condensed.

4 Focus Coil – Magnet 1 Cooling / Testing Cryo cooler operation check LN2 Cooling (Cryo cooler overnight) Coolers Operating LN2 Cooling Cold Mass temps Radiation Shield 2 nd Stage Head

5 Focus Coil – Magnet 1 Cooling / Testing Cooling to Full Gas input from bottle pack – 1.05 bar Level sensor operating but not able to communicate with the control system. System cooling over the Christmas period By the New year the level was reported as being 100%. Pressure feedback heater systems operating with a regulation of 0.005 bar. Level reduced to 80% using the pressure regulation heaters and the exhaust poppet valve. Shows a system that can be adjusted without issue Pressure regulated at 1.14 bar with a gas bottle backup set to 1.05 bar. System safety backup In the event of heater failure gas bottle would over pressurize. Easily demonstrated zero boil off conditions. Control system and monitoring works well.

6 Focus Coil – Magnet 1 Cooling / Testing Solenoid mode powering Powering in Solenoid mode starts 11 th January. Quench induced at 50A to test the dump resistors and quench system Some minor issues casing false quenches, problems solved. Power supply controller Threshold of the voltage limit to low – raised for following tests to 500mV Solenoid mode operating current of 114A achieved, held and ramped down. Magnet cabling re-configured to operate in Flip mode (225A operating current) Scope logging applied to the Quench system to log the quench events Quenches occur at 145A, 152A, 158A, 161A and 162A Concern over the lack of significant jumps in current Tesla stop the testing until understanding of the quenches.

7 Focus Coil – Magnet 1 Cooling / Testing

8 Continued Flip mode powering Following discussions between the MICE AFC group and Tesla is was decided to continue with testing Investigation of the voltage at the splice joints is to be looked into, 1 splice / coil Ramping up to a nominal current ~ 100A Continue to quench Quench occurred at 170A A significant jump from the last interval Plans are to continue applying current, Tesla have ordered more LHe to be delivered this week.

9 Focus Coil – Magnet 2 Construction Progress Construction The outer vacuum vessel is being leak tested at the moment Background He rate is high so causing some problems in ascertaining the actual leak rate of the vessel. Tesla are confident of there being no leaks in the vessel Radiation shield is nearing completion. Problem found with a sensor so intervention needed MLI wrapping of the radiation shield will be completed this week Cold Mass Will be inserted and mounted onto the supports this week Alignment will be carried out with oversight by Jason Tarrant He Turret The second He turret is being constructed in parallel to the main magnet construction. It is intended to be shipped with the magnet Oversight and Schedule Andy Nichols is now making regular visits to discuss with Tesla the progress and scheduled delivery. The forecast is for the second Focus Coil to arrive at RAL in July

10 Magnetic Mitigation - Analysis Model created using Opera and analysis using Vector Fields Detailed model for Step IV (more analysis of Step VI) Beam line magnets included DSA area Current shielding walls As much ferrous material equipment as possible, size limitations. Consideration of the ISIS and MICE control rooms for personnel safety.

11 Magnetic Mitigation - Analysis

12 Magnetic Mitigation – West wall Cryo Cooler head compressors - Single and Two stage Analysis has shown that the original location under the south mezzanine is unsuitable Field in the area around 400 gauss Local shielding would be required for each compressor High cost ~ £15k per compressor Moving the compressors to the West wall Reduces the ambient field to around 5 gauss No need for local shield Furthest distance possible (compressor hoses) Suitable for Step IV and Step VI compressors Number of compressors at each stage Step IV – 15 Two stage and 4 Single stage Step V – 21 Two stage and 4 Single stage Step VI – 27 Two stage and 4 Single stage Limitations He pressure lines no longer than 30m Access and egress to the main door for magnet deliveries, and other large equipment Minimise the work to “changes” to the current hall arrangements

13 Magnetic Mitigation – West wall


15 Magnetic Mitigation – Rack room 2 and MLCR

16 Magnetic Mitigation – Partial Return Yoke Parallel analysis at Brook Haven mainly aimed at the Step VI configuration Analysis and design is in place for the Step IV configuration Work needs to be started to gain a cost and resourced schedule for the installation effort. Changes to South Mezzanine Positioning of the Tracker cryostats Sliding floor changes Trench cabling / infrastructure changes Logistics

17 RF – DL and Hall Operations First Amplifier has reached 1.2 MW. Changes to the contact arrangements have been made to reduce air –side electrical breakdown Re-testing has re-started but are having trouble with the crowbar high voltage breakdown safety device. Current test has reached 125KW RF Engineer is now in post and has started work for the current testing period Also working on the requirements for the TIARA grant Water system for the RF Amplifier Installation of the Amplifier, Coax and dummy load

18 Hall Operations Decay Solenoid The magnet has been cooled in preparation for EMR running The new quench protection system will be tested with the old power supply The new power supply is scheduled for delivery in July Spectrometer Solenoid Frames Both frames have been installed into position The upstream frame has been loaded (5 tonnes) and test survey alignment carried out Model of a solenoid manufacture to aid with the alignment of the tracker patch panel and other service arrangements. LN2 system A new Hydrogen transfer line has been designed and procured in preparation for installation for Step IV operations

Download ppt "MICE - UK Project Progress Roy Preece 30 th April 2013."

Similar presentations

Ads by Google